40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 10 16:52:04 2020, gautam, Update, BHD, Workable setup prepared 
    Reply  Tue Aug 11 23:20:29 2020, gautam, Update, BHD, Some first tests with air BHD setup PRMI_RFlock.pdfPRMI_RFlock_fullscale.pdf
       Reply  Fri Aug 21 15:27:00 2020, gautam, Update, BHD, Better BHD mode-matching IMG_7548.JPGBHD_MM.pngPRMI_DCPDs.pdf
          Reply  Wed Aug 26 00:52:55 2020, gautam, Update, BHD, BHD activities 
Message ID: 15513     Entry time: Mon Aug 10 16:52:04 2020     Reply to this: 15514
Author: gautam 
Type: Update 
Category: BHD 
Subject: Workable setup prepared 

All the details are in E2000436, and documents linked from there, I think an elog would be much too verbose. In summary, a workable setup consisting of

  • 2 DCPDs interfaced with the realtime CDS system. Note that because this circuit is single-ended, while the AA and ADC are differential receiving, there is an overall gain of 0.5. Explicitly, for the 300 ohm DC transimpedance, the conversion is ~350 cts/mW.
  • A local oscillator beam delivered via fiber that is mode-matched (roughly) with the IFO AS beam.
  • A PZT mounted mirror to control the homodyne phase. The PZT (S320) is an obsolete part and it's hard to find a datasheet for it, but if its specs are comparable to the more modern S330, the full stroke is 10 um, for a max applied voltage of 100 V DC, so 100nm/V. c.f. 200V for 3um full stroke of the Noliac.

was prepared.

Last night, I locked the PRMI with the carrier resonant, and convinced myself that the DCPD null stream was sensing the MICH degree of freedom (while it was locked on AS55_Q) with good SNR below ~60 Hz. Above ~60 Hz, in this configuration, the ADC noise was dominating, but by next week, I'll have a whitening board installed that will solve this particular issue. With the optical gain of MICH in this configuration, the ADC noise level was equivalent to ~500 nrad/rtHz of phase noise above ~60 Hz (plots later).

Now, I can think about how to commission this setup interferometrically.

ELOG V3.1.3-